# Aligning Agents like Large Language Models

### v1

@@ Abstract
Training agents to behave as desired in complex 3D environments from high-dimensional sensory information is challenging. Imitation learning from diverse human behavior provides a scalable approach for training an agent with a sensible behavioral prior, but such an agent may not perform the specific behaviors of interest when deployed. To address this issue, we draw an analogy between the undesirable behaviors of imitation learning agents and the unhelpful responses of unaligned large language models (LLMs). We then investigate how the procedure for aligning LLMs can be applied to aligning agents in a 3D environment from pixels. For our analysis, we utilize an academically illustrative part of a modern console game in which the human behavior distribution is multi-modal, but we want our agent to imitate a single mode of this behavior. We demonstrate that we can align our agent to consistently perform the desired mode, while providing insights and advice for successfully applying this approach to training agents. Project webpage at: https://adamjelley.github.io/aligning-agents-like-llms/.
Figure 1: Illustration of our approach for aligning a generally capable agent with a game designer’s intended goals or preferences. A general agent pre-trained to imitate a large diverse dataset of human gameplay provides a base agent which can be more easily fine-tuned to imitate a smaller task or demonstration dataset. This agent can then be further fine-tuned with reinforcement learning using a reward model learned from preferences to reliably achieve the target behavior with the desired style. This approach is analogous to the alignment procedure used for recent large language models.
@@ 1 Introduction
The optimal approach for training generally competent agents to act in complex 3D environments without carefully crafted reward functions is an open question. Many modern console games provide such 3D environments, where the state space, action space and temporal horizons are large enough that learning from scratch is usually infeasible, even with a clearly defined goal. In these environments, a natural approach is to leverage a large dataset of general human behavior to pre-train an agent with imitation learning. This provides an agent with a general understanding of player behavior, and there is evidence of generalization benefits from scaling up data and compute (Lee et al., 2022; Baker et al., 2022; Reed et al., 2022; Brohan et al., 2023; Bousmalis et al., 2023; SIMA-Team et al., 2024).
@@
However, such an agent will inevitably learn to imitate all behaviors found within the human gameplay, including undesirable behaviors of novice or malicious players, analogous to unhelpful or toxic responses of unaligned large language models (Ziegler et al., 2020). Additionally, game designers may have preferences for agents to act with a certain style or strategy, for which it may be difficult to codify a suitable reward function (Aytemiz et al., 2021; Devlin et al., 2021). There is a parallel between training useful agents and aligning large language models. In the same way that LLMs have been aligned to serve as customer service chatbots (Banerjee et al., 2023), search engines (Spatharioti et al., 2023) and code generation assistants (Chen et al., 2021b), we can imagine aligning a large imitation learning agent with various objectives within a game. For example, we may wish to train agents that can act as allies, opponents (Vinyals et al., 2019), non-player-characters (NPCs) (Alonso et al., 2020), or even for quality assurance during development (Bergdahl et al., 2021).
In this work, we take inspiration from the current procedure for training large language models (LLMs) and investigate applying this procedure to train agents. We consider a situation where the human behavior distribution is distinctly multi-modal over non-negligible time horizons (of the order of $\sim 10$ seconds), but we desire our agent to imitate a single mode of this distribution. This particular mode of behavior may be difficult to learn from scratch with reinforcement learning since there is no clear reward function, but an imitation learning agent pre-trained on general gameplay data would only sometimes sample the desired behavior mode. At the same time, it may be infeasible to obtain sufficient clean demonstrations for robust imitation learning. To isolate this problem, we focus on an academically illustrative part of a modern console game where players must navigate from a randomly selected spawn point to one of three jumppads. We find that a base imitation learning agent reaches all three jumppads in similar proportions to the human data. We then demonstrate that we can align this agent to consistently reach a single preferred jumppad, using synthetic preference labelling to train a reward model followed by online reinforcement learning using this reward model.
Our contributions include: 1) An analysis of the importance and potential difficulties of applying each stage of the current LLM training pipeline to agents, including unsupervised pre-training, supervised fine-tuning, preference modelling and online alignment, 2) Evidence of representation transfer between imitation learning and preference learning that highlights the benefits of using the base agent to initialise the reward model when learning preferences from pixels, and 3) The introduction of an additional training stage, preference fine-tuning, to substantially improve alignment efficiency. We perform the entire procedure from visual inputs with full controller action output, maintaining the generality of our approach to other game environments, and demonstrating that the modern LLM training paradigm can be successfully applied to aligning agents on a real 3D game.
2 Game Environment and Alignment Goal
2.1 Environment
@@
For this paper, we utilize the video game Bleeding Edge [^1] https://www.bleedingedge.com/en, which was launched in 2020 for Xbox One [^2] https://www.pcgamingwiki.com/wiki/Bleeding_Edge. It is a team-based 4v4 online multi-player video game. Players select from thirteen possible heroes, each with different abilities. The game is played with a third-person view, with the camera angle controlled by the player, so the environment is partially observable. Here we focus on a single map, called Skygarden. This map is spread over three islands each with multiple elevation levels, including a main island and two launch islands (one for each team).
2.2 Alignment Objective
@@
At the beginning of the game, players spawn at one of four nearby points on their team’s launch island. From the spawn location a player must navigate across the launch island to one of three jumppads that launch the player from their launch island onto the main island. Depending on the jumppad selected, the player will be launched onto different areas of the main island, so players may wish to take different jumppads during a match depending on the location of opposing team players. For the purpose of this work, we aim to train an agent to consistently navigate across the launch island to a single one of the three jumppads. Since this navigation task requires around 10 seconds to complete for an optimal agent, a random agent will rarely leave the spawn area. Additionally, without access to privileged information such as the agent location (we only provide visual input as below) and an externally shaped reward function, it would be difficult to train an agent to complete this task. On the other hand, an imitation learning agent will learn to reach all three of the jumppads, as in the human distribution. Therefore while the task is simple, it provides a clear motivation for alignment and a setting which we can quantitatively analyse.
@@
Figure 2: Screenshots of the agent at a spawn point (2(a)), heading towards the middle jumppad (2(b)), and looking back at the launch island after having launched from the right jumppad (2(c)).
2.3 Observation and Action Spaces
@@
To maintain the generality of our approach to any 3D environment, we take visual gameplay observations $\mathbf{o}_{t}\in\mathbb{R}^{H\times W\times 3}$ as input, sampled at 10Hz. We do not provide any privileged information, so the agent only has access to input information available to human players. The action space consists of 12 buttons and 2 joysticks. The left joystick controls the movement of the agent, while the right joystick controls the camera angle. We decompose each joystick into an $x$ and a $y$ component which are independently discretised into 11 buckets. This creates a $12+2+2=16$ dimensional discrete action space in total, although the joystick dimensions are of most relevance to our objective.
3 Implementation and Analysis
3.1 Training a Base Imitation Learning Model
We now follow the general procedure outlined in Figure 1 and Appendix A and describe our specific implementation of this procedure for training an agent aligned to reliably reach a preferred jumppad. We begin by performing imitation learning from visual observations to gamepad actions on general gameplay data to provide a behavioral prior with a general understanding of human gameplay.
@@ 3.1 Training a Base Imitation Learning Model
Dataset: For general pre-training, a dataset was extracted from recorded human gameplay, as described in Appendix B. This large unfiltered dataset consists of 71,940 individual player trajectories from 8788 matches recorded between 09-02-2020 and 10-19-2022, which amounts to 9875 hours (1.12 years) of individual gameplay. For the purposes of this work, we use an agent that was trained for roughly one epoch on this dataset. Considering the success of recent foundation models, we note that given our unified observation and action spaces, it may also be possible to train across games to obtain a base model, as explored in previous work (Reed et al., 2022; SIMA-Team et al., 2024).
@@
Architecture and Training: For the policy, we use a GPT-2 (Radford et al., 2019) causal transformer architecture with 103M parameters, similar to that used by VPT (Baker et al., 2022). Observations from the human gameplay $\mathbf{o}_{t}\in\mathbb{R}^{H\times W\times 3}$ are taken directly as input to a convolutional encoder to give observation embeddings $\mathbf{z}_{t}$. The transformer is trained with a context window of $H=32$ timesteps (corresponding to around 3s of gameplay given the 10Hz sampling). The context window is important since the game is partially observable: as the context window is increased, the agent is able to better capture the state of the environment and take more informed actions (i.e. with a more Markovian state) at the cost of computational complexity. The output corresponds to the 16 discrete action dimensions. The transformer and convolutional encoder are both trained end to end with a cross-entropy loss over all output action components to provide a policy $\pi(a_{t}|o_{t},...o_{t-H})$. Full architecture and training details are provided in Appendix C.
@@
Evaluation: Once trained, we ran our base pre-trained agent online in the game environment and recorded which jumppad was reached for 1000 episodes. To run the agent online, we initialized the Ninja character with an empty context buffer at one of the spawn points at random. We ran the agent until it reached one of the jumppads or timed out. The jumppad distribution reached is demonstrated in Figure 3 below.
Figure 3: Distribution of jumppads reached by the base imitation learning agent.
@@
We find that our base model reaches a jumppad $56\%$ of the time, and has a bias towards the middle jumppad, reflecting human behavior. Our base model therefore provides a reasonable behavioral prior that is much better than random exploration of the launch island (which has a $0\%$ success rate). However, the success rate could be improved, which is likely due to distribution shift between the offline data and the online environment. For example, the base model was trained on data containing all thirteen possible characters, while our online evaluation only uses the default Ninja character. Additionally, we initialize our agents online at the spawn point with an empty context buffer, while in training the agent will have access to context including a spawn animation and prior gameplay. While measures can be taken to avoid distribution shift, offline only learning is inherently challenging (Ostrovski et al., 2021) and some distribution shift is usually inevitable, as we discuss further in Appendix D.
3.2 Supervised Fine-Tuning on Task Relevant Dataset
Figure 4: Distribution of jumppads reached by the fine-tuned imitation learning agent.
Following the LLM pipeline, we now fine-tune our agent on a smaller curated dataset. Here we fine-tune on a high quality subset of the pre-training data. However, more generally this could consist of fine-tuning on demonstrations of desireable behavior by the game designer.
@@ 3.2 Supervised Fine-Tuning on Task Relevant Dataset
Dataset: We curated 300 successful trajectories (100 per jumppad) for fine-tuning, all of which involved the character Ninja and were filtered to contain only the first part of the trajectory until a jumppad was reached. We fine-tune until convergence as described in Appendix C.
Results: We evaluate our agent following the same procedure as before. We see in Figure 4 that the success rate of reaching a jumppad is now substantially higher ($89\%$), and that the agent more evenly reaches all three jumppads (corresponding to the balanced fine-tuning dataset). At this stage, it is natural to ask whether the general pre-training was beneficial or whether the model could have been trained from scratch on our task relevant dataset directly. We find that the agent without pre-training has a less diverse jumppad distribution, and a higher failure rate (around $20\%$ of trajectories fail to reach a jumppad compared to only around $10\%$ for the pre-trained agent). Interestingly, from inspecting agent behaviors we found that the pre-trained agent appeared to be more robust since it can often better ‘self-correct’ its trajectories if it drifts out-of-distribution of the fine-tuning dataset. Full details of our pre-training ablation and an example of this self-correction behavior are provided in Appendix E.1. These results corroborate similar findings for LLMs, in which pre-training transformer models has been shown to enable more effective fine-tuning, effectively increasing the size of the fine-tuning data compared to training from scratch (Hernandez et al., 2021).
We also perform a preliminary investigation of model scaling by training additional models from scratch on this smaller task-specific dataset in Appendix E.2. We find that smaller models have a higher failure rate, even trained only on this relatively small fine-tuning dataset. While larger models than we consider may perform even better (particularly when incorporating pre-training), this would increase the inference cost which has a substantial impact on online alignment speed (and the potential for the final agent to behave in real-time), so we proceed with the 103M parameter model described above.
3.3 Obtaining Preference Data on Online Rollouts

### v2

@@ Abstract
Training agents to act competently in complex 3D environments from high-dimensional visual information is challenging. Reinforcement learning is conventionally used to train such agents, but requires a carefully designed reward function, and is difficult to scale to obtain robust agents that generalize to new tasks. In contrast, Large Language Models (LLMs) demonstrate impressively general capabilities resulting from large-scale pre-training and post-training alignment, but struggle to act in complex environments. This position paper draws explicit analogies between decision-making agents and LLMs, and argues that agents should be trained like LLMs to achieve more general, robust, and aligned behaviors. We provide a proof-of-concept to demonstrate how the procedure for training LLMs can be used to train an agent in a 3D video game environment from pixels. We investigate the importance of each stage of the LLM training pipeline, while providing guidance and insights for successfully applying this approach to agents. Our paper provides an alternative perspective to contemporary LLM Agents on how recent progress in LLMs can be leveraged for decision-making agents, and we hope will illuminate a path towards developing more generally capable agents for video games and beyond. Project summary and videos: https://adamjelley.github.io/aligning-agents-like-llms.
Machine Learning, ICML
@@ 1 Introduction
The optimal approach for training agents to act in complex 3D environments without well-defined reward functions is an open question. Many modern video games provide such environments, where the state space, action space, and time-horizons are large enough that reinforcement learning (RL) from scratch is usually infeasible. However, even if possible, the resulting agent would not generalize to new objectives or environments. In contrast, large language models (LLMs) demonstrate impressively general capabilities resulting from large-scale pre-training, but struggle to act in complex low-level environments (paglieri_balrog_2024).
Figure 1: Illustration of our approach for training general agents in complex environments. Pre-training on a large, diverse dataset of interactions provides a base agent which can be more effectively fine-tuned with limited demonstrations. This agent can be further refined with RL, using a reward model or external reward function. This approach is analogous to the training of modern LLMs.
An analogous approach in the context of training agents is to leverage a large dataset of general human gameplay to pre-train an agent with imitation learning (IL). This provides an agent with a general understanding of player behavior, and there is evidence of generalization benefits from scaling up data and compute (lee_multi-game_2022; baker_video_2022; reed_generalist_2022; brohan_rt-2_2023; pearce_scaling_2024).
@@
However, such an agent will inevitably learn to imitate all behaviors found within the human gameplay, including undesirable behaviors of novice or malicious players, analogous to unhelpful or toxic responses of unaligned LLMs (ziegler_fine-tuning_2020). Additionally, game designers may want this general agent to perform a particular role, or act with a certain style or strategy (aytemiz_acting_2021; devlin_navigation_2021). There is a parallel between making such an agent useful and aligning LLMs. In the same way that general LLM base models have been aligned for various purposes (banerjee_benchmarking_2023; spatharioti_comparing_2023; chen_evaluating_2021), we can imagine aligning a general decision-making agent with various objectives or environments.
This position paper argues that decision-making agents should be trained in the same manner as large language models (LLMs) to obtain analogous generality and robustness to objectives and environments. Specifically, agents should be trained and refined over multiple stages, typically involving: 1) Unsupervised pre-training on general behavior or gameplay data to obtain a general behavioral prior, 2) Supervised fine-tuning (SFT) on higher quality demonstration data of the relevant task or environment, 3) Reinforcement learning (RL), potentially from human feedback (RLHF), to refine and improve the behavior or further align the agent with implicit behavior preferences. To provide evidence for our position, we consider an academically illustrative part of a modern video game where the human behavior distribution is distinctly multi-modal, but we desire our agent to imitate a single mode of this distribution. This particular mode of behavior is difficult to learn from scratch with RL from pixels due to the size of the exploration space, and there are insufficient clean demonstrations of this mode for robust imitation learning. We demonstrate that this behavior mode can be reliably achieved by following the modern LLM training pipeline, and illustrate the importance of each stage. We train our agent from high-dimensional visual input with full game controller action output, maintaining the generality of our approach to other game environments, and demonstrating that the modern LLM training paradigm can be successfully applied to training agents in complex 3D environments. Our proof of concept provides evidence that many of the recent findings in the training of LLMs do indeed transfer to training agents, and could enable the training of more general and robust decision-making agents.
2 Context and Background
Large Scale Imitation Learning: Scaling up imitation learning and offline reinforcement learning (levine_offline_2020) has recently gained popularity with the aim of demonstrating similar scaling laws to LLMs (kaplan_scaling_2020; hoffmann_training_2022) to obtain more general agents. For example, Decision Transformer (chen_decision_2021) proposed imitation learning with a transformer policy that can be conditioned on a desired return. Multi-Game Decision Transformers (lee_multi-game_2022) extended this approach to learn multiple game policies, and GATO (reed_generalist_2022) demonstrated that a single model could learn hundreds of diverse tasks. RT1 (brohan_rt-1_2022) shows similar scaling potential in robotics tasks, while RT2 (brohan_rt-2_2023) integrated vision-language models and RoboCat (bousmalis_robocat_2023) demonstrated promising adaptation to new tasks. SIMA (sima-team_scaling_2024) incorporates language conditioning to enable generalization to new video games. Recently, pearce_scaling_2024 have confirmed the hypothesized scaling laws for imitation learning. While these approaches show the potential for scaling imitation learning, we argue that these models are analogous to GPT-2 (radford_improving_2018) in the rise of modern LLMs and that additional alignment stages can similarly be used to specialize and improve these models to make them as useful as modern LLMs.
Fine-tuning with Reinforcement Learning: The use of imitation learning as pre-training for reinforcement learning was first investigated to improve the sample efficiency of deep RL algorithms by reducing the exploration space (hester_deep_2017; vecerik_leveraging_2018). AlphaGo (silver_mastering_2016) and subsequently AlphaStar (vinyals_grandmaster_2019) demonstrated that large-scale imitation learning could provide a strong behavioral prior for environments like StarCraft where RL from scratch is infeasible. VPT (baker_video_2022) extended this paradigm to internet-scale Minecraft videos using an inverse dynamics model to enable imitation learning, before RL fine-tuning on task-specific rewards. While these works demonstrate the potential of fine-tuning large imitation models, they use hard-coded reward functions to maximize performance rather than align an agent’s behavior with new objectives or subjective preferences.
Reinforcement Learning from Human Feedback (RLHF): Training agents with human preferences has a long history, notably including knox_tamer_2008 and bennett_netflix_2009, which has been reviewed by wirth_survey_2017 and zhang_recent_2021. The popular modern formulation of RLHF for deep learning was proposed by christiano_deep_2017. This idea was extended by ibarz_reward_2018 to include imitation learning as pre-training to improve the efficiency of reward modeling. PEBBLE (lee_pebble_2021) instead uses unsupervised pre-training to increase the diversity of initial behaviors. abramson_improving_2022 and milani_towards_2023 are closest to our position in using RLHF to refine the performance of large imitation agents in 3D simulated worlds, but do not make an explicit analogy or investigate the full LLM training pipeline.
Large Language and Vision Models: radford_improving_2018 demonstrated that pre-training with an unsupervised generative task (next token prediction) on a large diverse corpus of text, followed by fine-tuning on a specific task, led to performance benefits compared to fine-tuning alone. stiennon_learning_2020, following ziegler_fine-tuning_2020, then popularized the use of RLHF to further fine-tune these models to align their responses with human preferences. This procedure led to InstructGPT (ouyang_training_2022), Chat-GPT (openai_introducing_2022), and GPT-4 (openai_gpt-4_2023) as well as a wide range of other popular language models, such as Claude (bai_training_2022), Llama (touvron_llama_2023) and Gemini (geminiteam2024geminifamilyhighlycapable). Generative pre-training has been shown to have equivalent benefits for vision tasks (chen_generative_2020), and RLHF has similarly been used to refine image generation and vision-language models (VLMs) (xu_imagereward_2023; esser_scaling_2024). More recently, RL with pre-trained LLMs has been shown to improve reasoning capabilities in verifiable domains where external rewards can be defined (openai_openai_2024; deepseek-ai_deepseek-r1_2025). While components of the LLM training pipeline have been applied both individually and in combination to decision-making agents, in this paper we argue that the full pipeline (outlined in Appendix A) should be more widely and explicitly applied to agents to achieve similarly general agent models. We now turn to providing evidence to support this position with a concrete proof-of-concept.
3 Proof of Concept
3.1 Environment and Alignment Goal
@@
As a proof of concept, we use the AAA video game Bleeding Edge111www.bleedingedge.com/en, which was launched in 2020 for Xbox One222www.pcgamingwiki.com/wiki/Bleeding_Edge. It is a team-based 4v4 online multi-player video game. The game is played with a third-person view, with the camera angle controlled by the player, so the environment is partially observable. Here, we focus on a single map, called Skygarden. This map is spread over three islands, including a main island and two launch islands (one for each team).
@@
At the beginning of a game, players spawn at one of four points on their team’s launch island. From the spawn point, a player must navigate across the launch island to one of three jumppads that launch the player onto the main island. Depending on the jumppad selected, the player will be launched onto a different area of the main island, so players often take different jumppads. For the purpose of this work, we aim to train an agent to navigate across the launch island to a single one of the three jumppads. Since this navigation task requires around 10 seconds to complete for an optimal agent, without access to privileged information such as the agent location to shape a reward function, it would be extremely difficult to train an agent to complete this task with RL from visual input alone, as it would rarely leave the spawn area. On the other hand, an imitation learning agent will match the human distribution to reach all three of the jumppads. Therefore while the task is simple, it provides a clear motivation for alignment and a setting which we can quantitatively analyze to illustrate our position.
3.2 Implementation Details
@@
Figure 2: Screenshots of the Ninja agent at a spawn point and heading towards the middle jumppad of the launch island.
@@
Observation and Action Spaces: To maintain the generality of our approach to any 3D environment, we take visual gameplay observations $\mathbf{o}_{t}\in\mathbb{R}^{H\times W\times 3}$ as input, sampled at 10Hz. We do not provide any privileged information, so the agent only has access to visual information, as shown in Figure 2. The action space consists of 12 buttons and 2 joysticks. The left joystick controls movement, while the right controls the camera angle. We decompose each joystick into an $x$ and a $y$ component which are independently discretized into 11 buckets to enable better modeling of multi-modal behavior distributions (compared to directly using continuous values), which is essential for our setting.
@@
Dataset: For general pre-training, a dataset was extracted from recorded human gameplay, as described in Appendix B. This large unfiltered dataset consists of 71,940 individual player trajectories from 8788 matches recorded between 09-02-2020 and 10-19-2022, which amounts to 9875 hours (1.12 years) of individual gameplay. For the purposes of this work, we use a base agent that was trained for roughly one epoch on this dataset. Considering the success of recent foundation models, we note that given our unified observation and action spaces, it may be possible to instead train across games to obtain this base model, as explored in previous work (reed_generalist_2022; sima-team_scaling_2024).
@@
Architecture and Training: For the policy, we use a GPT-2 (radford_language_2019) causal transformer architecture with 103M parameters, similar to that used by VPT (baker_video_2022). Observations from the human gameplay $\mathbf{o}_{t}\in\mathbb{R}^{H\times W\times 3}$ are taken directly as input to a convolutional encoder to give observation embeddings $\mathbf{z}_{t}$. The transformer is trained with a context window of $H=32$ timesteps (around 3s of gameplay given the 10Hz sampling). The context window is important since the game is partially observable: longer context provides the agent with a more Markovian state, but at the cost of quadratic computational complexity. We note that developments in extending the context length of transformers, such as modified attention mechanisms (beltagy_longformer_2020), or even Retrieval-Augmented Generation (RAG) (lewis_retrieval-augmented_2021), could also be investigated here to help address partial observability. The transformer and convolutional encoder are both trained end to end with a cross-entropy loss over all 16 discrete action components to provide a policy $\pi(a_{t}|o_{t},a_{t-1}...o_{t-H})$. While we only consider prior observation and action context for simplicity in this proof-of-concept, text embeddings $e$ corresponding to game or task instructions (likely to be an integral part of a general agent) could also be straightforwardly included as conditioning here, providing an instruction-conditioned policy $\pi(a_{t}|o_{t},a_{t-1}...o_{t-H};e_{t})$ similar to SIMA (sima-team_scaling_2024). Full architecture and training details are provided in Appendix C.
@@
Evaluation: To evaluate our models, we run each agent online in the game environment and record which jumppad is reached over 1000 episodes. To run an agent online, we initialized the Ninja character with an empty context buffer at one of the spawn points at random. We run the agent until it reaches one of the jumppads or times out.
3.3 Does large scale pre-training provide generalization benefits in the context of agents?
Figure 3: Distribution of jumppads reached by the base agent.
@@ 3.3 Does large scale pre-training provide generalization benefits in the context of agents?
We begin by evaluating our base model, which achieves the jumppad distribution shown in Figure 3. We see that our agent fails to reach a jumppad $44\%$ of the time, and has a bias towards the middle jumppad, reflecting human behavior. Our base model therefore provides a reasonable behavioral prior, and is much better than a randomly initialized agent (which has a $0\%$ success rate). However, the success rate could be improved, which is partly due to distribution shift between the offline data and the online environment. For example, the base model was trained on data containing all thirteen possible characters, while our online evaluation only uses the default Ninja character. Additionally, we initialize our agents online at the spawn point with an empty context buffer, while in training the agent will have access to context including a spawn animation and prior gameplay. While measures can be taken to avoid distribution shift, offline-only learning is inherently challenging (ostrovski_difficulty_2021) as discussed in Appendix D, demonstrating the need for some online behavior refinement (see Section 3.6).
@@
Following the LLM pipeline, we now fine-tune our agent on a smaller curated dataset. We curated 300 successful trajectories (100 per jumppad) for fine-tuning, all of which involved the character Ninja and were filtered to contain only the first part of the trajectory until a jumppad was reached. We fine-tune until convergence as described in Appendix C. More generally, this stage could consist of fine-tuning on any smaller, high-quality demonstration data. We see in Figure 4 that the success rate of reaching a jumppad is now substantially higher (with only an $11\%$ failure rate), and that the agent more evenly reaches all three jumppads (corresponding to the balanced fine-tuning dataset).
Figure 4: Distribution of jumppads reached by the base agent after supervised fine-tuning (SFT) on a task-specific dataset.
At this stage, it is natural to ask whether the general pre-training was beneficial or whether the model could have been trained from scratch on our task-relevant dataset directly. After all, if clean demonstrations of the desired behavior are available, why train the agent on messy, less-relevant data? To investigate this, we train an equivalent architecture model from scratch until convergence on the fine-tuning dataset only (see Appendix C for details). The evaluation of this model is shown below in Figure 5.
Figure 5: Distribution of jumppads reached by an equivalent agent trained from scratch on the task-specific dataset used for supervised fine-tuning.
Perhaps surprisingly, we find that the agent without pre-training has a less diverse jumppad distribution, and a higher failure rate (around $20\%$, almost twice that of the pre-trained agent). This finding seems to go against conventional wisdom that only clean, relevant demonstrations should be used from the available data. Instead, these results corroborate similar findings for LLMs, in which pre-training on diverse data has been shown to enable more effective fine-tuning, effectively increasing the size of the fine-tuning data compared to training from scratch (hernandez_scaling_2021).
To understand the reason behind this difference, we reviewed videos of the two agents navigating towards the jumppads. We noticed that failures would often arise when an agent accumulates action error over its trajectory, a common challenge with imitation learning (ross_efficient_2010), causing it to miss the intended jumppad as shown in Figure 6. Since the fine-tuning dataset consists of only a small number of trajectories that hit the jumppad directly, the agent trained from scratch (without pre-training) has never seen that observation before, so therefore chooses the modal action (which is to move forwards) as it is now out of distribution. This causes it to run into the wall continuously until it times out, leading to failure (Fig. 6, left). On the other hand, the pre-trained agent would often hit the wall, and then turn around and hit the jumppad from behind (Fig. 6, right). This is because the more diverse pre-training data provides a common-sense behavior prior, allowing the agent to return to the fine-tuning distribution. These behaviors are demonstrated in the videos on our project webpage, and provide anecdotal but intuitive evidence for the benefits of unsupervised pre-training in the context of agents.
Figure 6: Screenshots of the agent after having missed the intended jumppad. Left: Agent without pre-training continues into wall. Right: Agent with pre-training turns around to return to jumppad.
We also perform a preliminary investigation of model scaling in Appendix E, by training additional models with an equivalent transformer architecture from scratch on this smaller task-specific dataset. We find that smaller models have a higher failure rate, providing evidence that larger model sizes lead to improved performance in our imitation learning task, even when trained only on this relatively small fine-tuning dataset. These findings are in agreement with more comprehensive recent work by pearce_scaling_2024. Furthermore, we see early indications that larger models also align more efficiently with reinforcement learning.
In summary, we find that large-scale pre-training
provides general benefits to agents, analogous to LLMs.
3.4 Preference Modeling